OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and parallel to y = 2x + 4. y = -2x + 7 y = 2x - 3 y = 2x - 7 y = -2x + 3

OpenStudy (anonymous):

If two lines are parallel, they have the same gradient. So start off by looking at the gradient of the line y=2x +4. You then need to work out the y-intercept of your new line. To do this, use the general equation of a straight line, y=mx +c. m is the gradient that you just found. c is the y-intercept. To find c, then substitute the y and x values from the point (1,-5) and rearrange for c. :)

OpenStudy (anonymous):

I'm still confused d:

OpenStudy (anonymous):

Not a problem, on which part exactly? :P

OpenStudy (anonymous):

Do you have the gradient?

OpenStudy (anonymous):

I'm confused about it all!

OpenStudy (anonymous):

The general equation of a straight line is y=mx +c. m is the gradient. c is the y-intercept. The equation you have is y=2x+4, which is of the same form, y=mx+c. And so, comparing the two equations, the line y=2x+4 must have a gradient of 2. Is this bit okay?

OpenStudy (anonymous):

Hmmm, let me try it

OpenStudy (anonymous):

Would it be B?

OpenStudy (anonymous):

I think the answer is y = 2x - 7. What did you do?

OpenStudy (anonymous):

I guessed :P You said the gradient has to be two so that narrowed it down to b and c :P

OpenStudy (anonymous):

Haha! Don't do that in any exams ;) Firstly, do you understand why the gradient of y=2x+4 is 2?

OpenStudy (anonymous):

Yes. M is the move right? And 4 is y intercept.

OpenStudy (anonymous):

Perfect. Next, do you understand that if two lines are parallel, they have the same gradient? Or 'move' :P

OpenStudy (anonymous):

I'm starting to get it..

OpenStudy (anonymous):

Because if one line sloped more than the other one, then they would cross at some point. So we need them never to cross if they are parallel, and so they should have the same slope.

OpenStudy (anonymous):

Oh okay :P

OpenStudy (anonymous):

Can you help me with a few more, Maybe i will understand better with a different question.

OpenStudy (anonymous):

I can try :) Do you know how to then work out why the y=intercept is -7 on the one we just looked at?

OpenStudy (anonymous):

Not so much :P I'm new at this

OpenStudy (anonymous):

Which of the following is the slope between the points (0, -10) and (-4, 0)? A:2/5 B:-2/5 C: -5/2 D: 5/2

OpenStudy (anonymous):

So, you know the gradient of your new line is 2. But you don't know its y intercept. So, you have: y=2x+c You know that this line goes through the point (1,-5). So, substitute the values into the equation. y=2x+c becomes -5=2*1 +c so, c=-5-2=-7

OpenStudy (anonymous):

That was for the previous question.

OpenStudy (anonymous):

For the next one, draw a graph and plot the two points. Draw a line between them. Are you aware that the gradient of the line is equal to (change in the y coordinate)/(change in the x coordinate) ?

OpenStudy (anonymous):

I'm sorry lol i'm trying to answer other questions. Lol

OpenStudy (anonymous):

And i'm not aware.. Like i said.. I am new

OpenStudy (anonymous):

So the gradient of a line is equal to: (change in the y coordinate)/(change in the x coordinate) So, going from the point (0,-10) to (-4,0), The y coordinate increases by 10, so 10 is your change in y. Your a coordinate decreases by 4, so -4 is your change in x. You then substitute these values into the equation to get the gradient. It does help to plot the points and the line, because then you can label your change in y and your change in x :)